OpenStudy (anonymous):

Find the LCM of 21x^2 + 63x and 7x^2 + 49x + 84 =

OpenStudy (dumbcow):

\[21x^{2}+63x = 21x(x+3)\] \[7x^{2}+49x+84 = 7(x^{2}+7x+12) = 7(x+4)(x+3)\] LCM is product of factors without double counting common factors LCM of 7 and 21 is 21 \[LCM \rightarrow 21x(x+4)(x+3)\]